Mathematics 243 – Algebraic Structures Definitions Quiz 2 – September 15, 2006

Your Name:

A) (5) Let ∗ be a binary operation on a set A. What does it mean to say that ∗ is associative.

B) (5) Let ∗ be the binary operation on Z defined by x ∗ y = x + y − 7. Is there an identity element for ∗? If there is, say what it is and why it works; if not say why not.